What is the pKa of an acid that has a Ka value of 7.38 x 10-7?

a.

6.13

b.

7.87

c.

6.63

d.

7
Answer #1

a) 6.13

pKa = - logKa

Given

Ka =  7.38 x 10-7

pKa = - log [ 7.38 x 10-7]

pKa = 7 - log 7.38 = 7- 0.868 = 6.13